A 75 kg object moving at 2.5 m/s is slowed by a 25 N resistance force over a distance of 9m. What is the final velocity of the object?
Answer #1

finding the acceleration first by a=F/m=25/75=0.333m/s^2

now using v^2=u^2+2as

v=sqrt(2.5^2-2*0.333*9)= 0.505 m/s